### v1

# Restless Bandit Problem with Rewards Generated by a Linear Gaussian Dynamical System

Abstract
The stochastic multi-armed bandit problem studies decision-making under uncertainty. In the problem, the learner interacts with an environment by choosing an action at each round, where a round is an instance of an interaction. In response, the environment reveals a reward, which is sampled from a stochastic process, to the learner. The goal of the learner is to maximize cumulative reward. A specific variation of the stochastic multi-armed bandit problem is the restless bandit, where the reward for each action is sampled from a Markov chain. The restless bandit with a discrete state-space is a well-studied problem, but to the best of our knowledge, not many results exist for the continuous state-space version which has many applications such as hyperparameter optimization. In this work, we tackle the restless bandit with continuous state-space by assuming the rewards are the inner product of an action vector and a state vector generated by a linear Gaussian dynamical system. To predict the reward for each action, we propose a method that takes a linear combination of previously observed rewards for predicting each action’s next reward. We show that, regardless of the sequence of previous actions chosen, the reward sampled for any previously chosen action can be used for predicting another action’s future reward, i.e. the reward sampled for action 1 at round $t-1$ can be used for predicting the reward for action $2$ at round $t$. This is accomplished by designing a modified Kalman filter with a matrix representation that can be learned for reward prediction. Numerical evaluations are carried out on a set of linear Gaussian dynamical systems.
1 Introduction
The Stochastic Multi-Armed Bandit (SMAB) problem provides a rigorous framework for studying decision-making under uncertainty. The problem consists of the interaction between a learner and an environment for a set number of rounds. For each round, the learner chooses an action and in response the environment reveals a reward, which is sampled from a stochastic process, to the learner. The goal of the learner is to maximize cumulative reward. In the non-stationary case of the SMAB, the distributions of the reward for each action can change each round. A key result in the area is [1] where it assumes that the cumulative changes in the reward distributions are bounded by a known constant.
A more specific variation of the non-stationary SMAB are rewards are generated by $s$-step autoregressive models, i.e. an action’s sampled reward $X_{t}$ is a linear combination of rewards $X_{t-s},\dots,X_{t-1}$ where $s$ is the autoregressive model order. Two key results that have tackled this SMAB environment are [2], [3], and [4]. [2] studied the performance of a number of algorithms for rewards generated by Brownian motion. In [3], the authors consider when the rewards for each action is generated by a known 1-step autoregressive process. In [4], they address SMAB environments modeled as a unknown 1-step autoregressive or a known $s$-step autoregressive. A key application of autoregressive models is presented in [5], where the work tunes the hyperparameters, such as the gradient descent’s learning rate, during the training process of a reinforcement learning neural network. Finally, another perspective to $s$-step autoregressive models is [6] where rewards $X_{t}$ and a context $\theta_{t}$ are generated by a Linear Gaussian Dynamical System (LGDS). The authors prove that a linear combination of previously observed contexts $\theta_{t-s},\dots,\theta_{t-1}$ can be used to predict the reward $X_{t}$, a perspective similar to the environments considered in [3] and [4].
Figure 1: In this work, we illustrate that for the proposed LGDS, the linear combination of any sequence of observed rewards can be used to predict each action’s reward. For example, if we take the previously observed rewards $X_{t-3}$, $X_{t-2}$, $X_{t-1}$ based on the chosen actions $A_{t-3}=3$, $A_{t-2}=2$, $A_{t-1}=1$, we can predict the reward for actions $a=1,2,3$ using the linear combination of previously observed rewards. The linear combination is based on an estimated vector $\hat{G}_{a|\mathbf{c}}$, $a\in\{1,2,3\}$.
Our work proposes a discrete-time restless bandit with continuous state-space by assuming the state and rewards are generated by a LGDS. This paper extends the results in [6] where now the context is absent. The contributions of our paper are as follows.
Our Contributions:
  • We introduce a discrete-time, continuous state-space, restless bandit where the rewards are generated by a LGDS in Section 2.
  • We prove that we can predict the reward for each action by using a linear combination of observed rewards. The coefficients for the linear combination are from the identified modified Kalman filter matrix representation. (See Section 3)
  • We provide a proof of the error bound of the reward prediction for the identified modified Kalman filter. The idea is inspired by [7] for identifying the Kalman filter, where now we assume that the measurements of the LGDS, a linear combination of the system’s state, can change each round.
  • Using the proved error bound of the reward prediction, we propose the algorithm Optimistic Sensor Selector (OSS). The algorithm chooses the action that maximizes the sum of the reward prediction and its error. (See Section 4)
  • For numerical results in Section 5, we apply OSS to a parameterized LGDS to illustrate its numerical performance. Here, we compare OSS to Upper Confidence Bound (UCB) algorithm [8] and Sliding Window UCB (SW-UCB) [9] algorithm, two well-known SMAB algorithms, and for which LGDS OSS performs best.
Related Work
An environment considered in non-stationary SMAB is the restless bandit in [10]. In the restless bandit, the reward for each action is the function of a state that is generated by a Markov chain. Whenever the learner chooses an action, the learner observes a Markov chain’s state and a reward. This paper focuses on the case when the transition matrix of the Markov chain is unknown. Previous results in the discrete state-space Markov chain that use an approach similar to UCB are [11, 12, 13, 14, 15]. [16] uses Thompson sampling, i.e. sampling parameters based on a priori distribution of Markov chain, for action selection. We avoid comparisons with these previous results since the states of the Markov chain are discrete, whereas the results presented in this paper focus on when the states are continuous. This allows us to tackle a different set of application, such as hyperparameter optimization for reinforcement learning neural networks, e.g. [5].
2 Problem Formulation
For this paper, the learner will interact with an environment modeled as a LGDS. We will consider the following LGDS where the rewards are the inner product of an action vector and the LGDS’s state:
$\begin{cases}z_{t+1}&=\Gamma z_{t}+\xi_{t},~{}~{}z_{0}\sim\mathcal{N}\left(\hat{z}_{0},P_{0}\right)\\ X_{t}&=\left\langle c_{A_{t}},z_{t}\right\rangle+\eta_{t}\end{cases},$ (1)
where $z_{t}\in\mathbb{R}^{d}$ is the state of the system and $X_{t}\in\mathbb{R}$ is the reward. The process noise $\xi_{t}\in\mathbb{R}^{d}$ and measurement noise $\eta_{t}\in\mathbb{R}$ are independent normally distributed random variables, i.e. $\xi_{t}\sim\mathcal{N}(0,Q)$ and $\eta_{t}\sim\mathcal{N}(0,\sigma)$. The vector $c_{A_{t}}\in\mathcal{A}=\{c_{a}\in\mathbb{R}^{d\times 1}\mid\left\|c_{a}\right\|_{2}\leq B_{c},a\in[k]\}$ where $B_{c}$ is known and $a\in[k]\triangleq\{1,2,\dots,k\}$ is the indexed action. Using similar notation as [17], actions that are chosen at round $t$ are denoted as $c_{A_{t}}\in\mathcal{A}$; otherwise, the action is denoted as $c_{a}\in\mathcal{A}$. We make the following assumptions on system (1).
Assumption 1.
The state matrix $\Gamma$ is marginally stable, i.e. $\rho\left(\Gamma\right)\leq 1$.
Assumption 2.
The vectors and matrices in system (1) are unknown along with $Q$, $\sigma$, and $d$. However, number of actions $k$ is known.
Assumption 3.
The matrix pair $\left(\Gamma,Q^{1/2}\right)$ is controllable. The pair $\left(\Gamma,c_{a}^{\top}\right)$ is detectable for every vector $c_{a}\in\mathcal{A}$.
The goal of the learner is to maximize the cumulative reward over a horizon $n>0$, i.e. $\sum_{t=1}^{n}X_{t}$. The horizon length $n>0$ may be unknown. To provide analysis on the performance of any proposed algorithm for maximizing cumulative reward in (1), regret is analyzed which is defined to be
$R_{n}\triangleq\sum_{t=1}^{n}\mathbb{E}\left[X_{t}^{*}-X_{t}\right],$ (2)
where $X_{t}^{*}$ is the highest possible reward that can be sampled at round $t$. In the next section, we discuss a reward predictor for the LGDS (1).
3 Predicting the Reward of the LGDS
This section reviews the optimal $1$-step predictor of the rewards, in the mean-squared error sense, generated by LGDS (1): the Kalman filter. According to Assumption 2, the matrices of the LGDS (1) are unknown, implying that the Kalman filter needs to be identified. However, direct identification of the Kalman filter for the LGDS (1) is intractable. Therefore, we propose a modified Kalman filter to identify. Imposing the assumptions posed in the previous section, we prove that prediction error of the modified Kalman filter is lower than the variance of the reward $X_{t}$, making it a sufficiently useful predictor of the reward. The added benefit of the modified Kalman filter is that it is tractable to identify.
The Kalman filter uses the previous observations $X_{1},\dots,X_{t}$ to compute an estimate of the state $z_{t}$ as $\hat{z}_{t}\triangleq\mathbb{E}\left[z_{t}\mid\mathcal{F}_{t-1}\right]$ where $\mathcal{F}_{t-1}$ is the sigma algebra generated by the rewards $X_{1},\dots,X_{t-1}$,
$\begin{cases}\hat{z}_{t+1}&=\Gamma\hat{z}_{t}+\Gamma K_{t}\left(X_{t}-\left\langle c_{A_{t}},\hat{z}_{t}\right\rangle\right),~{}~{}P_{t+1}=g\left(P_{t},c_{A_{t}}\right)\\ K_{t}&=P_{t}c_{A_{t}}\left(c_{A_{t}}^{\top}P_{t}c_{A_{t}}+\sigma\right)^{-1}\\ \hat{X}_{t}&=\left\langle c_{A_{t}},\hat{z}_{t}\right\rangle\end{cases},$ (3)
and $g\left(P,c\right)$ is defined to be the following Riccati equation [18]
$g\left(P,c\right)\triangleq\Gamma P\Gamma^{\top}+Q-\Gamma Pc\left(c^{\top}Pc+\sigma\right)^{-1}c^{\top}P\Gamma^{\top}.$ (4)
We impose the following assumption for the LGDS’s (1) initial state $z_{0}\sim\mathcal{N}(\hat{z}_{0},P_{0})$ and the Kalman filter’s (3) initial error covariance matrix $P_{0}$:
Assumption 4.
The initial state $z_{0}\in\mathbb{R}^{d}$ of the LGDS (1) is sampled from a normal distribution with mean $\hat{z}_{0}=\lim_{\tau\rightarrow\infty}\Gamma^{\tau}\hat{z}$, where $\hat{z}\in\mathbb{R}^{d}$ is an arbitrary vector, and covariance matrix $P_{0}$. We assume for $P_{0}$ that $P_{0}=P_{\overline{a}}$, where $P_{\overline{a}}$ is the steady-state error covariance matrix, $P_{\overline{a}}=g\left(P_{\overline{a}},c_{\overline{a}}\right)$, $c_{\overline{a}}\in\mathcal{A}$.
Remark 1.
Assumption 4 states that LGDS (1) is in steady-state and the Kalman filter’s (3) error covariance matrix is bounded. This is a reasonable assumption as the Kalman filter covariance matrix $P_{t}$ converges exponentially to the steady state covariance matrix $P_{\overline{a}}$ as $t$ increases if action $c_{\overline{a}}\in\mathcal{A}$ is consistently chosen. In addition, a similar assumption has been made in [19], [20], and [7]. Finally, it will be proven in Lemma 1 that there exists an action $c_{\overline{a}}\in\mathcal{A}$ such that $P_{\overline{a}}\succeq P_{t}$ if $P_{\overline{a}}=P_{0}$.
As mentioned earlier, the parameters of LGDS (1) are unknown due to Assumption 2. Therefore, we propose to learn the Kalman filter (3) for reward prediction. However, since the Kalman filter matrices $P_{t}$ and $K_{t}$ change constantly, it is intractable to identify the Kalman filter. Therefore, we prove that there exists a modified Kalman filter that has a bounded reward prediction error regardless of the choices $c_{A_{t}}\in\mathcal{A}$ that is tractable to identify. For proving Theorem 1, we first provide Lemma 1 for the bound on the Kalman filter error covariance matrix $P_{t}$.
Lemma 1.
Let $P_{a}$, $a\in[k]$ be the steady state solution of the Kalman filter for each action $c_{a}\in\mathcal{A}$, $P_{a}=g\left(P_{a},c_{a}\right)$, where $g\left(P_{a},c_{a}\right)$ is defined in (4). Define $P_{\overline{a}}\succeq 0$ to be the steady-state error covariance matrix of the Kalman filter (3) associated with action $c_{\overline{a}}\in\mathcal{A}$ such that $P_{\overline{a}}\succeq P_{a}$ for every action $a\in[k]$. By imposing Assumptions 1, 3, and 4, the LGDS (1), then $P_{\overline{a}}\succeq P_{t}$ for any $t=1,2,\dots,n$.
Below is Theorem 1 which proves the existence of a modified Kalman filter with a bounded prediction error. Proof for Theorem 1 can be found in Appendix B.
Theorem 1.
We define the following modified Kalman filter
$\begin{cases}\hat{z}_{t+1}^{\prime}&=\Gamma\hat{z}_{t}^{\prime}+\Gamma L_{A_{t}}\left(X_{t}-\left\langle c_{A_{t}},\hat{z}_{t}^{\prime}\right\rangle\right)\\ X_{t}&=\left\langle c_{A_{t}},\hat{z}_{t}^{\prime}\right\rangle+\gamma_{A_{t}}\end{cases},~{}~{}L_{A_{t}}\triangleq P_{\overline{a}}c_{A_{t}}\left(c_{A_{t}}^{\top}P_{\overline{a}}c_{A_{t}}+\sigma\right)^{-1},$ (5)
where $\gamma_{A_{t}}\triangleq X_{t}-\left\langle c_{A_{t}},\hat{z}_{t}^{\prime}\right\rangle\sim\mathcal{N}\left(0,c_{A_{t}}^{\top}P_{t}^{\prime}c_{A_{t}}+\sigma\right)$ and $P_{t}^{\prime}\triangleq\mathbb{E}\left[\left(z_{t}-\hat{z}_{t}^{\prime}\right)\left(z_{t}-\hat{z}_{t}^{\prime}\right)^{\top}\mid\mathcal{F}_{t-1}\right]$. It is proven for the modified Kalman filter (5) that 1) the matrix $\Gamma-\Gamma L_{A_{t}}c_{A_{t}}^{\top}$ is stable and 2) the variance of the residual $\mbox{Var}\left(\gamma_{A_{t}}\right)$ is bounded.
The key takeaway for Theorem 1 is that there exists a modified Kalman filter (5) that is easier to identify in comparison to the Kalman filter (3) at the expense of a higher prediction error $\mbox{Var}\left(\gamma_{a}\right)\geq c_{a}^{\top}P_{t}c_{a}+\sigma^{2}$. This is because the modified Kalman filter has only a finite number of gain matrices $L_{A_{t}}$ and a static covariance matrix $P_{\overline{a}}$. In addition, the variance of the prediction error $\mbox{Var}\left(\gamma_{a}\right)$ has an upper-bound.
3.1 Learning the Modified Kalman filter
Using Theorem 1 and inspired by the results presented in [7], we will learn the modified Kalman filter. Let parameter $s>0$ denote how far in the past the learner will look. We define the tuple $\mathbf{c}\triangleq\begin{pmatrix}c_{A_{t-s}}&\dots&c_{A_{t-1}}\end{pmatrix}$ as the sequence of actions chosen by the learner from rounds $t-s$ to $t-1$. The reward $X_{t}=\left\langle c_{a},z_{t}\right\rangle+\eta_{t}$ for action $a\in[k]$ can be expressed as a linear combination of rewards $X_{t-s},\dots,X_{t-1}$ generated by the tuple $\mathbf{c}$ using the matrices defined in the modified Kalman filter (5):
[…]
Therefore, let there be defined the vectors $G_{c_{a}|\mathbf{c}}$, $c_{a}\in\mathcal{A}$, and $\Xi\left(\mathbf{c}\right)$ to express the reward $X_{t}=\left\langle c_{a},z_{t}\right\rangle+\eta_{t}$:
$\Rightarrow X_{t}=G_{c_{a}\mid\mathbf{c}}^{\top}\Xi_{t}\left(\mathbf{c}\right)+\beta_{a}+\gamma_{a},$ (7)
$G_{c_{a}\mid\mathbf{c}}$ $\triangleq\begin{bmatrix}c_{a}^{\top}\left(\Gamma-\Gamma L_{A_{t-1}}\right)\cdots\left(\Gamma-\Gamma L_{A_{t-s+1}}\right)\Gamma L_{A_{t-s}}&\dots&c_{a}^{\top}\Gamma L_{A_{t-1}}\end{bmatrix}\in\mathbb{R}^{s\times 1}$
$\Xi_{t}\left(\mathbf{c}\right)$ $\triangleq\begin{bmatrix}X_{t-s}&\dots&X_{t-1}\end{bmatrix}^{\top}\in\mathbb{R}^{s\times 1}$
$\beta_{a}$ $\triangleq c_{a}^{\top}\left(\Gamma-\Gamma L_{A_{t-1}}c_{A_{t-1}}^{\top}\right)\cdots\left(\Gamma-\Gamma L_{A_{t-s}}c_{A_{t-s}}^{\top}\right)\hat{z}_{t-s}^{\prime}\in\mathbb{R}.$
Based on equation (7), we can express the reward $X_{t}=\left\langle c_{a},z_{t}\right\rangle+\eta_{t}$ for each action $c_{a}\in\mathcal{A}$ using $G_{c_{a}|\mathbf{c}}$, $c_{a}\in\mathcal{A}$, and $\Xi\left(\mathbf{c}\right)$ with the following linear model:
$\begin{pmatrix}\left\langle c_{1},z_{t}\right\rangle+\eta_{t}\\ \left\langle c_{2},z_{t}\right\rangle+\eta_{t}\\ \vdots\\ \left\langle c_{k},z_{t}\right\rangle+\eta_{t}\\ \end{pmatrix}=\begin{pmatrix}G_{c_{1}\mid\mathbf{c}}^{\top}\\ G_{c_{2}\mid\mathbf{c}}^{\top}\\ \vdots\\ G_{c_{k}\mid\mathbf{c}}^{\top}\end{pmatrix}\Xi_{t}\left(\mathbf{c}\right)+\begin{pmatrix}\beta_{1}\\ \beta_{2}\\ \vdots\\ \beta_{k}\end{pmatrix}+\begin{pmatrix}\gamma_{1}\\ \gamma_{2}\\ \vdots\\ \gamma_{k}\end{pmatrix}.$ (8)
The linear model (8) proves that we only need to identify $k^{s+1}$ vectors $G_{c_{a}\mid\mathbf{c}}$. Therefore, we can 1) identify $G_{c_{a}\mid\mathbf{c}}$ for each action $c_{a}\in\mathcal{A}$ and 2) predict the reward $X_{t}$ using inner product of the identified $G_{c_{a}\mid\mathbf{c}}$ and sequence of rewards $X_{t-s},\dots,X_{t-1}$.
Remark 2.
In the linear model (8), there is a parameter $s>0$ which is the number of previous rewards $X_{t-s},\dots,X_{t-1}$ used for predicting the reward $X_{t}$. Parameter $s>0$ impacts the magnitude of the term $\beta_{a}$ (which decreases exponentially as $s$ increases) and the number of linear models $G_{c_{a}\mid\mathbf{c}}$ to identify (which increases exponentially as $s$ increases).
The following are assumed about $G_{c_{a}\mid\mathbf{c}}$, $\gamma_{a}$, and $c_{a}\in\mathcal{A}$:
Assumption 5.
There exists a known upper bound $B_{G}$ such that $\|G_{c_{a}\mid\mathbf{c}}\|_{2}\leq B_{G}$ for all $a\in[k]$ which is a common assumption to use in SMAB problems [21].
Assumption 6.
There exists a known constant $B_{R}>0$ such that for any round $t>0$, we have:
$\sqrt{\mbox{tr}\left(Z_{t}\right)}\leq B_{R},~{}Z_{t}\triangleq\mathbb{E}\left[z_{t}z_{t}^{\top}\right],~{}\mbox{Var}\left(\gamma_{a}\right)\leq c_{a}^{\top}P_{\overline{a}}c_{a}+\sigma_{\eta}\leq B_{R}^{2}\mbox{ for }c_{a}\in\mathcal{A},$
where $Z_{t}$ (which has the iteration $Z_{t+1}=\Gamma Z_{t}\Gamma^{\top}+Q$) is the covariance of the LGDS’s (1) state $z_{t}$.
To learn $G_{c_{a}\mid\mathbf{c}}$, assume that at time points $\mathcal{T}_{c_{a}\mid\mathbf{c}}=\left\{t_{1},\dots,t_{N_{a}}\right\}$ ($N_{a}$ is the number of times action $c_{a}\in\mathcal{A}$ is chosen) the following tuple sequence $\begin{pmatrix}c_{A_{t_{i}-s}},\dots,c_{A_{t_{i}-1}}\end{pmatrix}=\mathbf{c}\in\mathcal{A}^{s}\mbox{ for }t_{i}\in\left\{t_{1},\dots,t_{N_{a}}\right\}$ and action $c_{A_{t}}=c_{a}\in\mathcal{A}$ are chosen. We have the following linear model
$\mathbf{X}_{\mathcal{T}_{c_{a}\mid\mathbf{c}}}=G_{c_{a}\mid\mathbf{c}}^{\top}\mathbf{Z}_{\mathcal{T}_{c_{a}\mid\mathbf{c}}}+\mathbf{B}_{\mathcal{T}_{c_{a}\mid\mathbf{c}}}+\mathbf{E}_{\mathcal{T}_{c_{a}\mid\mathbf{c}}},$ (9)
$\mathbf{X}_{\mathcal{T}_{c_{a}\mid\mathbf{c}}}$ $\triangleq\begin{bmatrix}X_{t_{1}}&\dots&X_{t_{N_{a}}}\end{bmatrix}\in\mathbb{R}^{1\times N_{a}},$ $\mathbf{Z}_{\mathcal{T}_{c_{a}\mid\mathbf{c}}}$ $\triangleq\begin{bmatrix}\Xi_{t_{1}}\left(\mathbf{c}\right)&\dots&\Xi_{t_{N_{a}}}\left(\mathbf{c}\right)\end{bmatrix}\in\mathbb{R}^{s\times N_{a}},$
$\mathbf{B}_{\mathcal{T}_{c_{a}\mid\mathbf{c}}}$ $\triangleq\begin{bmatrix}\beta_{A_{t_{1}}}&\dots&\beta_{A_{t_{N_{a}}}}\end{bmatrix}\in\mathbb{R}^{1\times N_{a}},$ $\mathbf{E}_{\mathcal{T}_{c_{a}\mid\mathbf{c}}}$ $\triangleq\begin{bmatrix}\gamma_{A_{t_{1}}}&\dots&\gamma_{A_{t_{N_{a}}}}\end{bmatrix}\in\mathbb{R}^{1\times N_{a}}.$
The least squares estimate of $G_{c_{a}\mid\mathbf{c}}$ in (9) is
$\hat{G}_{c_{a}\mid\mathbf{c}}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)$ $=\mathbf{X}_{\mathcal{T}_{c_{a}\mid\mathbf{c}}}\mathbf{Z}_{\mathcal{T}_{c_{a}\mid\mathbf{c}}}^{\top}V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)^{-1}$ (10)
$V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)$ $\triangleq\lambda I_{s}+\mathbf{Z}_{\mathcal{T}_{c_{a}\mid\mathbf{c}}}\mathbf{Z}_{\mathcal{T}_{c_{a}\mid\mathbf{c}}}^{\top}=\lambda I_{s}+\sum_{i=1}^{N_{a}}\Xi_{t_{i}}\left(\mathbf{c}\right)\Xi_{t_{i}}\left(\mathbf{c}\right)^{\top},$ (11)
where $\lambda>0$ is a regularization term. Since there are $k^{s}$ codes $\mathbf{c}\in\mathcal{A}^{s}$, then there are $k^{s+1}$ vectors $G_{c_{a}\mid\mathbf{c}}$ to learn.
4 Optimistic Sensor Selector LGDS Restless Bandit Problem
The section above provided a predictor, the modified Kalman filter, for the rewards generated by the LGDS (1). It also provided a methodology for identifying the predictor. Now that the reward can be predicted using an identified modified Kalman filter, we discuss how to use the predictor in Algorithm 1, Optimistic Sensor Selector (OSS). For each round $t$ in OSS, the learner will choose actions based on the following optimization problem
$\underset{a\in[k]}{\arg\max}~{}\hat{G}_{c_{a}\mid\mathbf{c}}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)^{\top}\Xi_{t}\left(\mathbf{c}\right)+\left(e_{c_{a}\mid\mathbf{c}}\left(\delta\right)+b_{c_{a}\mid\mathbf{c}}\left(\delta\right)\right)\sqrt{\Xi_{t}\left(\mathbf{c}\right)^{\top}V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)^{-1}\Xi_{t}\left(\mathbf{c}\right)},$ (12)
where with a probability of at least $(1-\delta)^{2}$, the following inequality is satisfied (see Theorem 4 in Appendix C.2):
$\hat{G}_{c_{a}\mid\mathbf{c}}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)^{\top}\Xi_{t}\left(\mathbf{c}\right)-G_{c_{a}\mid\mathbf{c}}^{\top}\Xi_{t}\left(\mathbf{c}\right)\leq\left(e_{c_{a}\mid\mathbf{c}}\left(\delta\right)+b_{c_{a}\mid\mathbf{c}}\left(\delta\right)\right)\sqrt{\Xi_{t}\left(\mathbf{c}\right)^{\top}V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)^{-1}\Xi_{t}\left(\mathbf{c}\right)}.$ (13)
The terms $e_{c_{a}\mid\mathbf{c}}\left(\delta\right)$ and $b_{c_{a}\mid\mathbf{c}}\left(\delta\right)$ are defined as
$e_{c_{a}\mid\mathbf{c}}\left(\delta\right)$ $\triangleq\sqrt{2B_{R}^{2}\log\left(\frac{1}{\delta}\frac{\det(V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right))^{1/2}}{\det(\lambda I)^{1/2}}\right)}$ (14)
$b_{c_{a}\mid\mathbf{c}}\left(\delta\right)$ $\triangleq\sqrt{N_{a}}\frac{B_{c}B_{R}}{\delta}\sqrt{\mbox{tr}\left(I-\lambda V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)^{-1}\right)}+\lambda\sqrt{\mbox{tr}\left(V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)^{-1}\right)}B_{G}.$ (15)
Reward prediction uncertainty (13) of action $c_{a}\in\mathcal{A}$ is impacted directly $V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)$ which is a sum of $N_{a}$ (number of times action $c_{a}\in\mathcal{A}$ is chosen) positive semi-definite matrices. Therefore, choosing an action frequently (large $N_{a}$) will lower the reward prediction uncertainty. The rationale behind optimization problem (12) is to balance choosing the action with the highest reward versus the action with the most uncertainty. We summarize below which term is defined to be in (12) within Algorithm 1 which consists of an Exploitation term (which action the learner expects to return the highest reward) and an Exploration term (how much should the learner explore an action).
  • Exploitation term: $\hat{G}_{c_{a}\mid\mathbf{c}}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)^{\top}\Xi_{t}\left(\mathbf{c}\right)$
  • Exploration term: $\left(e_{c_{a}\mid\mathbf{c}}\left(\delta\right)+b_{c_{a}\mid\mathbf{c}}\left(\delta\right)\right)\sqrt{\Xi_{t}\left(\mathbf{c}\right)^{\top}V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)^{-1}\Xi_{t}\left(\mathbf{c}\right)}$
Parameter $\delta$ is a failure rate of the bound in (13) where $\delta$ values closer to 0 computes a larger bound (13). Parameter $s$ is the number of previously observed rewards $X_{t-\tau}$ ($\tau=1,\dots,s$) used for predicting the next reward. The number of models to learn increases exponentially as $s$ increases. Finally, $\lambda$ is a regularization parameter to ensure that (11) is invertible.
Input: $\delta\in(0,1)$, $\lambda>0$, $s\in\mathbb{N}$, $B_{c},B_{R},B_{G}\in\mathbb{R}^{+}$
// Initialization
for $\mathbf{c}\in\left\{\begin{pmatrix}c_{a_{1}}&c_{a_{2}}&\dots&c_{a_{s}}\end{pmatrix}\in\mathcal{A}^{s}\right\}$ do
       for $a\in[k]$ do
             $\mathcal{T}_{c_{a}\mid\mathbf{c}}\leftarrow\{\},~{}~{}V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)\leftarrow\lambda I_{s},~{}~{}\hat{G}_{c_{a}\mid\mathbf{c}}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)\leftarrow\mathbf{0}_{s\times 1}$
             $\left(e_{c_{a}\mid\mathbf{c}}\left(\delta\right),b_{c_{a}\mid\mathbf{c}}\left(\delta\right)\right)\leftarrow 1/\epsilon$ where $\epsilon$ small
       end for
end for
// Learner interaction with LGDS
for $t=1,2,\dots,n$ do
       if  $t\geq s$  then
             // Action selection
             $A_{t}\leftarrow\underset{a\in\{1,2,\dots,k\}}{\arg\max}\hat{G}_{c_{a}\mid\mathbf{c}}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)^{\top}\Xi_{t}\left(\mathbf{c}\right)+\left(e_{c_{a}\mid\mathbf{c}}\left(\delta\right)+b_{c_{a}\mid\mathbf{c}}\left(\delta\right)\right)\sqrt{\Xi_{t}\left(\mathbf{c}\right)^{\top}V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)^{-1}\Xi_{t}\left(\mathbf{c}\right)}$
             Sample $X_{t}$ from (1)
             $\mathbf{c}\leftarrow\begin{pmatrix}c_{A_{t-s}}&\dots&c_{A_{t-1}}\end{pmatrix}$
             $\mathcal{T}_{c_{A_{t}}\mid\mathbf{c}}\leftarrow\mathcal{T}_{c_{A_{t}}\mid\mathbf{c}}\cup\{t\}$
             // Update estimates
             $V_{A_{t}}\left(\mathcal{T}_{c_{A_{t}}\mid\mathbf{c}}\right)\leftarrow V_{A_{t}}\left(\mathcal{T}_{c_{A_{t}}\mid\mathbf{c}}\right)+\Xi_{t}\left(\mathbf{c}\right)\Xi_{t}\left(\mathbf{c}\right)^{\top}$
             $\hat{G}_{c_{A_{t}}\mid\mathbf{c}}\left(\mathcal{T}_{c_{A_{t}}\mid\mathbf{c}}\right)\leftarrow\hat{G}_{c_{A_{t}}\mid\mathbf{c}}\left(\mathcal{T}_{c_{A_{t}}\mid\mathbf{c}}\right)^{\top}+X_{t}\Xi_{t}\left(\mathbf{c}\right)^{\top}V_{A_{t}}\left(\mathcal{T}_{c_{A_{t}}\mid\mathbf{c}}\right)^{-1}$
             // Update bounds
             Set $e_{c_{A_{t}}\mid\mathbf{c}}\left(\delta\right)$ and $b_{c_{A_{t}}\mid\mathbf{c}}\left(\delta\right)$ based on (14) and (15), respectively.
      else
             $A_{t}\leftarrow\mbox{Sample uniformly }a\sim[k]$
             Sample $X_{t}$ from (1)
       end if
end for
Algorithm 1 Optimistic Sensor Selector (OSS)
4.1 Regret Performance
Algorithm 1, OSS, has the following upper bound on regret (2). Proof is in Appendix C.3.
Theorem 2.
Using Algorithm 1, regret (2) satisfies the following inequality with a probability of at least $(1-\delta)^{4}$:
$R_{n}\leq\sum_{a=1}^{k}2nB_{c}^{2}B_{R}^{2}\left(1-\left(1-\delta\right)^{4}\left(1-\exp\left(\frac{-4B\left(\delta,c_{a^{*}}\mid\mathbf{c}\right)^{2}}{2\Delta G_{c_{a}\mid\mathbf{c}}^{\top}\Sigma_{\Xi_{t}\left(\mathbf{c}\right)}\Delta G_{c_{a}\mid\mathbf{c}}}\right)\right)\right),$ (16)
where $\Delta G_{c_{a}\mid\mathbf{c}}\triangleq G_{c_{a^{*}}\mid\mathbf{c}}-G_{c_{a}\mid\mathbf{c}}$, $\Sigma_{\Xi_{t}\left(\mathbf{c}\right)}\triangleq\mathbb{E}\left[\Xi_{t}\left(\mathbf{c}\right)\Xi_{t}\left(\mathbf{c}\right)^{\top}\right]$, and $B\left(\delta,c_{a^{*}}\mid\mathbf{c}\right)$ is a bound on uncertainty (13).
Theorem 2 proves that regret increases at worst linearly, i.e. $\mathcal{O}\left(n\right)$. First, the regret bound (16) decreases as the difference $\Delta G_{c_{a}\mid\mathbf{c}}$ or the reward covariance $\Sigma_{\Xi_{t}\left(\mathbf{c}\right)}$ increase. Second, the regret bound (16) decreases as uncertainty (13) decreases. The bound for $B\left(\delta,c_{a_{t}^{*}}\mid\mathbf{c}\right)$ can be found in Appendix C.2.
Remark 3.
Although it is common in stationary bandits to see sublinear regret, linear regret may be the best possible rate of regret for non-stationary bandits (this is proven in [1]).
5 Numerical Results
For numerical results, we generated rewards $X_{t}\in\mathbb{R}$ for each action $\{c_{1},c_{2}\}$ from the following LGDS:
$\begin{cases}z_{t+1}&=\begin{pmatrix}0.9R\left(\theta\right)&I_{2}\\ \mathbf{0}_{2\times 2}&0.9R\left(\theta\right)\end{pmatrix}z_{t}+\xi_{t}\\ X_{t}&=\left\langle c_{A_{t}},z_{t}\right\rangle+\eta_{t}\end{cases},~{}\begin{cases}R\left(\theta\right)&\triangleq\begin{pmatrix}\cos\theta&\sin\theta\\ -\sin\theta&\cos\theta\end{pmatrix}\\ c_{1}&\triangleq\begin{pmatrix}10&0&0&0\end{pmatrix}\\ c_{2}&\triangleq\begin{pmatrix}0&10&0&0\end{pmatrix}\end{cases},$ (17)
where the process noise $\xi_{t}\sim\mathcal{N}\left(0,I_{4}\right)$ and the measurement noise $\eta_{t}\sim\mathcal{N}\left(0,1\right)$ are sampled from standard normal distributions. The LGDS (17) is proposed to study how the magnitude of the error covariance matrix $P_{\overline{a}}$ impacts performance of OSS, where $P_{\overline{a}}$ is directly impacted by the parameter $\theta\in[0,2\pi]$. Prior to the learner’s interaction with the LGDS (17), $10^{4}$ time steps are computed of the LGDS (17) to set the system to a steady state. After, the $10^{4}$ time steps, the length of the interaction between the environment and the learner is $n=10^{4}$ rounds. Regret (2) is used to provide a metric of performance. Parameter $s$ in algorithm 1, OSS, is set to $1$ in the top left plot. For comparison, we consider UCB [8], SW-UCB [9], and a learner that selects a random action each round (this learner is denoted as Random). We use UCB as an comparison since the eigenvalues of the LGDS (17) state matrix $\Gamma\in\mathbb{R}^{4\times 4}$ is Schur, implying that the reward distributions are stationary. SW-UCB is also used as a comparison since the reward is still generated by a dynamical system. Finally, Random is used as baseline for worst performance.
Figure 2: Comparison algorithm’s regret normalized with respect to OSS’s regret. A positive percent implies that OSS has a lower regret than the compared algorithm.
In the top left plot of Figure 2, regret for UCB (red), Sliding UCB (green) and Random (blue) is normalized with respect to OSS’s regret for each $\theta\in[0,2\pi]$. The middle plot of Figure 2 is the minimum eigenvalue of the Observability Gramian $\mathcal{O}$ [22] for both actions $c_{a}\in\{c_{1},c_{2}\}$, which is the solution of the Lyapunov equation $\mathcal{O}=\Gamma^{\top}\mathcal{O}\Gamma+c_{a}c_{a}^{\top}$. The bottom plot of Figure 2 is the real part of the eigenvalue of the state matrix $\Gamma$. In the white regions, all the comparison algorithms outperform OSS. Based on the plot in the middle, it appears that the low Observability Gramian minimum eigenvalue and a positive real part of the state matrix’s eigenvalue is the cause. For the blue regions, no algorithm outperforms Random, implying that the rewards are too noisy to estimate/predict for the compared algorithms. Finally, the gray regions is approximately where OSS performs the best, providing at least a 10% improvement for each of the algorithms mid-region. Based on the bottom 2 plots, this increase in performance is from the high observability and an eigenvalue with a negative real part for the state matrix. High observability lowers the magnitude of the error covariance matrix $P_{\overline{a}}$, which leads to a lower regret bound of OSS. In addition, an eigenvalue with a negative real part for the state matrix leads to rapid switching of the optimal action, making it difficult for UCB to adapt. For the plot on the far right, this is the relative performance of OSS for each parameter $s=1,2,3$ when the LGDS system (17) parameter set to approximately $85\pi/$ 8 (approximately where we see the largest improvement in performance of OSS in the top left plot). Therefore, it appears that as $s$ increases to $s=2,3$, regret performance of OSS decreases. In addition, regret for parameter $s=3$ is similar to regret of parameter $s=2$ as parameter $s=3$ may be identifying a more precise predictor but has more parameters to identify, which can cause a longer period of exploration.
6 Conclusion
We have presented an algorithm for addressing a variation of the restless bandit with a continuous state-space. The rewards generated by this restless bandit variation is a LGDS. Based on the formulation, we propose to learn a representation of the modified Kalman filter to predict the rewards for each action. We have shown that regardless of the sequence of actions chosen, the learned representation of the modified Kalman filter converges. It is then proven what strategy should be used given the bound on regret, leading to an optimism-based strategy.
In this work, we have not considered how the sequence of actions $c_{A_{t-s}},\dots,c_{A_{t-1}}$ impact prediction error and how to choose parameter $s$, i.e. how far the learner looks into the past. First, the perturbation added for exploration only considers error of the model and not the sequence of actions impact on the error of the prediction. In other words, the chosen sequence of actions are myopic. This is similar to what is studied in sensor selection problems where the sequence of chosen sensors impact the error of the Kalman filter state estimate. Therefore, future work will focus on the action sequence impact on the reward prediction. Next, an important parameter in OSS is the window size $s$, which is the number of rewards in the past that are used for predicting the reward for each action. In OSS, this is a parameter to set prior to the interaction with the environment. However, questions we care to ask is how to automate the process of choosing $s$.
Appendix A Proof of Lemma 1
Proof.
Let $P_{\overline{a}}$ be the steady-state solution of the Kalman filter (3) such that
$P_{\overline{a}}\succeq P_{a}\mbox{ for all }c_{a}\in\mathcal{A}\mbox{ where }P_{a}\mbox{ solves }P_{a}=g\left(P_{a},c_{a}\right).$
Based on [23], if $P_{\overline{a}}\succeq P$ for any arbitrary $P\succeq 0$, then based on definition of $g\left(\cdot,\cdot\right)$ in (3):
$g\left(P_{\overline{a}},c_{a}\right)\succeq g\left(P,c_{a}\right),$
for any $c_{a}\in\mathcal{A}$ that is detectable with the LGDS (1). Since the LGDS (1) is detectable for any $c_{a}\in\mathcal{A}$, then the Kalman filter’s difference equation in (3) converges exponentially to its steady-state solution [24]. In addition, $P_{\overline{a}}\succeq P_{a}$ for any $c_{a}\in\mathcal{A}$. Therefore, $P_{\overline{a}}\succeq g\left(P_{\overline{a}},c_{a}\right)$ for any $c_{a}\in\mathcal{A}$. Let there be the modified Kalman filter (5) and the error covariance matrix $P_{t}^{\prime}\triangleq\mathbb{E}\left[\left(z_{t}-\hat{z}_{t}^{\prime}\right)\left(z_{t}-\hat{z}_{t}^{\prime}\right)^{\top}\mid\mathcal{F}_{t-1}\right]$. The iteration of $P_{t}^{\prime}$ is
$P_{t+1}^{\prime}=\left(\Gamma-\Gamma L_{a_{t}}c_{A_{t}}^{\top}\right)P_{t}^{\prime}\left(\Gamma-\Gamma L_{a_{t}}c_{A_{t}}^{\top}\right)^{\top}+Q+\sigma^{2}\Gamma L_{a_{t}}L_{a_{t}}^{\top}\Gamma^{\top}$ (18)
where it is clear from arguments above that $P_{\overline{a}}\succeq P_{t}^{\prime}$ and $P_{\overline{a}}\succeq P_{t+1}^{\prime}$. ∎
Appendix B Proof of Theorem 1
Proof.
Using the fact from Lemma 1 that
$g\left(P_{\overline{a}},c_{a}\right)\succeq g\left(P_{t}^{\prime},c_{a}\right),$
if $P_{\overline{a}}\succeq P_{t}^{\prime}$ for any $c_{a}\in\mathcal{A}$ and
$\left(\Gamma-\Gamma L_{A_{t}}c_{A_{t}}^{\top}\right)P_{\overline{a}}\left(\Gamma-\Gamma L_{A_{t}}c_{A_{t}}^{\top}\right)^{\top}+Q+\sigma^{2}\Gamma L_{A_{t}}L_{A_{t}}^{\top}\Gamma^{\top}=g\left(P_{\overline{a}},c_{a}\right),$ (19)
then (18) has a steady-state solution for any $c_{A_{t}}\in\mathcal{A}$. Finally, based on Lyapunov method for (19), the matrix $\Gamma-\Gamma L_{a_{t}}c_{A_{t}}^{\top}$ is stable. In addition, there exists an action $c_{A_{t}}\in\mathcal{A}$ such that the error $\gamma_{A_{t}}\sim\mathcal{N}\left(0,c_{A_{t}}^{\top}P_{t}^{\prime}c_{A_{t}}+\sigma\right)$ has the following upper-bound
$c_{A_{t}}^{\top}P_{t}^{\prime}c_{A_{t}}+\sigma\leq c_{A_{t}}^{\top}P_{\overline{a}}c_{A_{t}}+\sigma.$
Appendix C Rationale behind OSS
To develop an algorithm for LGDS restless bandits (1), we propose the following methodology. First, we will propose that actions should be chosen based on the following optimization problem
$\underset{a\in[k]}{\arg\max}~{}\hat{G}_{c_{a}\mid\mathbf{c}}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)^{\top}\Xi_{t}\left(\mathbf{c}\right)+u_{a},$ (20)
where $\hat{G}_{c_{a}\mid\mathbf{c}}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)^{\top}\Xi_{t}\left(\mathbf{c}\right)$ is an Exploitation term (which action the learner predicts to have the highest reward) and an arbitrary term $u_{a}$ that is an Exploration term (how much should the learner continue exploring this action). We first prove the regret bound as a function of $u_{a}$. We then design $u_{a}$ such that the regret bound is minimized.
In Subsection C.1, the model error bound of $\hat{G}_{c_{a}\mid\mathbf{c}}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)$ is proven in Theorem 3. To prove Theorem 3, we provide Lemma 2. In Subsection C.2, prediction error $\hat{G}_{c_{a}\mid\mathbf{c}}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)^{\top}\Xi_{t}\left(\mathbf{c}\right)$ is proven. Theorem 4 proves the bound of the following difference:
$\hat{G}_{c_{a}\mid\mathbf{c}}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)^{\top}\Xi_{t}\left(\mathbf{c}\right)-G_{c_{a}\mid\mathbf{c}}^{\top}\Xi_{t}\left(\mathbf{c}\right).$ (21)
Lemma 3 is then provided to prove the rate of convergence for prediction error that is derived in Theorem 4. Finally, Subsection C.3 provides proofs of Algorithm 1 regret bound. We first prove a regret bound if actions are chosen based on optimization problem (20). Based on the derived bound, we propose what value $u_{a}$ should be set to in (20).
C.1 Model Error
The following lemma is provided which will be used for proving Theorem 3.
Lemma 2.
Let $\hat{G}_{c_{a}\mid\mathbf{c}}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)$ and $G_{c_{a}\mid\mathbf{c}}$ be as in (10) and (8), respectively. Then, the matrix of the bias terms $\mathbf{B}_{\mathcal{T}_{c_{A_{t}}\mid\mathbf{c}}}$ satisfies the following inequality with a probability of at least $1-\delta$ (the failure probability value $\delta$ is assumed to be in the open interval $(0,1)$):
$\left\|\mathbf{Z}_{\mathcal{T}_{c_{A_{t}}\mid\mathbf{c}}}\mathbf{B}_{\mathcal{T}_{c_{A_{t}}\mid\mathbf{c}}}^{\top}\right\|_{V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)^{-1}}\leq\sqrt{N_{a}}\frac{B_{c}B_{R}}{\delta}\sqrt{\mbox{tr}\left(I-\lambda V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)^{-1}\right)}.$ (22)
Proof.
According to (9), $\left\|\mathbf{Z}_{\mathcal{T}_{c_{A_{t}}\mid\mathbf{c}}}\mathbf{B}_{\mathcal{T}_{c_{A_{t}}\mid\mathbf{c}}}^{\top}\right\|_{V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)^{-1}}$ can be expressed as
$\left\|\mathbf{Z}_{\mathcal{T}_{c_{A_{t}}\mid\mathbf{c}}}\mathbf{B}_{\mathcal{T}_{c_{A_{t}}\mid\mathbf{c}}}^{\top}\right\|_{V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)^{-1}}=\left\|\sum_{t\in\mathcal{T}_{c_{a}\mid\mathbf{c}}}\Xi_{t}\left(\mathbf{c}\right)\beta_{A_{t}}\right\|_{V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)^{-1}}.$ (23)
Based on the triangle and Cauchy-Schwarz inequalities, (23) has the following bound
$\left\|\mathbf{Z}_{\mathcal{T}_{c_{A_{t}}\mid\mathbf{c}}}\mathbf{B}_{\mathcal{T}_{c_{A_{t}}\mid\mathbf{c}}}^{\top}\right\|_{V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)^{-1}}$ $=\left\|\Xi_{t_{1}}\left(\mathbf{c}\right)^{\top}\beta_{a;t_{1}}V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)^{-1/2}+\dots+\Xi_{t_{N_{a}}}\left(\mathbf{c}\right)^{\top}\beta_{a;t_{N_{a}}}V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)^{-1/2}\right\|_{2}$
$\overset{(*)}{\leq}\left\|\Xi_{t_{1}}\left(\mathbf{c}\right)^{\top}\beta_{a;t_{1}}V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)^{-1/2}\right\|_{2}+\dots$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}+\left\|\Xi_{t_{N_{a}}}\left(\mathbf{c}\right)^{\top}\beta_{a;t_{N_{a}}}V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)^{-1/2}\right\|_{2}$
$\overset{(**)}{\leq}\left|\beta_{a;t_{1}}\right|\left\|\Xi_{t_{1}}\left(\mathbf{c}\right)^{\top}V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)^{-1/2}\right\|_{2}+\dots$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}+\left|\beta_{a;t_{N_{a}}}\right|\left\|\Xi_{t_{N_{a}}}\left(\mathbf{c}\right)^{\top}V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)^{-1/2}\right\|_{2},$
$\Rightarrow\left\|\mathbf{Z}_{\mathcal{T}_{c_{A_{t}}\mid\mathbf{c}}}\mathbf{B}_{\mathcal{T}_{c_{A_{t}}\mid\mathbf{c}}}^{\top}\right\|_{V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)^{-1}}\leq\sum_{i=1}^{N_{a}}\left|\beta_{a;t_{i}}\right|\left\|\Xi_{t_{i}}\left(\mathbf{c}\right)^{\top}V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)^{-1/2}\right\|_{2},$ (24)
where ()(*)( ∗ ) uses the triangle inequality and ()(**)( ∗ ∗ ) uses the Cauchy-Schwarz inequality. The term $\left|\beta_{A_{t}}\right|$, $t\in\mathcal{T}_{c_{a}\mid\mathbf{c}}$ has the following bound using Cauchy-Schwarz inequality
$\left|\beta_{A_{t}}\right|$ $=\left|\left\langle c_{a},\left(\Gamma-\Gamma L_{a_{t-1}}c_{a_{t-1}}^{\top}\right)\cdots\left(\Gamma-\Gamma L_{a_{t-s}}c_{a_{t-s}}^{\top}\right)\hat{z}_{t-s}\right\rangle\right|$
$\leq\left\|c_{a}\right\|_{2}\left\|\left(\Gamma-\Gamma L_{a_{t-1}}c_{a_{t-1}}^{\top}\right)\cdots\left(\Gamma-\Gamma L_{a_{t-s}}c_{a_{t-s}}^{\top}\right)\hat{z}_{t-s}\right\|_{2}.$ (25)
We can upper-bound the right side of (25) with $\left\|\hat{z}_{t}\right\|_{2}$. First, define $\hat{Z}_{t}\triangleq\mathbb{E}\left[\hat{z}_{t}\hat{z}_{t}^{\top}\mid\mathcal{F}_{t-1}\right]$ to get the following two iteration
$\hat{Z}_{t-s+1}^{[1]}$ $=\left(\Gamma-\Gamma L_{a_{t-s}}c_{a_{t-s}}^{\top}\right)\hat{Z}_{t-s}^{[1]}\left(\Gamma-\Gamma L_{a_{t-s}}c_{a_{t-s}}^{\top}\right)^{\top}+\mathbb{E}\left[X_{t-s}^{2}\right]\Gamma L_{a_{t-s}}L_{a_{t-s}}^{\top}\Gamma^{\top}$
$\hat{Z}_{t-s+1}^{[2]}$ $=\left(\Gamma-\Gamma L_{a_{t-s}}c_{a_{t-s}}^{\top}\right)\hat{Z}_{t-s}^{[2]}\left(\Gamma-\Gamma L_{a_{t-s}}c_{a_{t-s}}^{\top}\right)^{\top}$
where $\sqrt{\mbox{tr}\left(\hat{Z}_{t}^{[1]}\right)}$ and $\sqrt{\mbox{tr}\left(\hat{Z}_{t}^{[2]}\right)}$ are equal to
$\sqrt{\mbox{tr}\left(\hat{Z}_{t}^{[1]}\right)}$ $=\left\|\hat{z}_{t}\right\|_{2}$ (26)
$\sqrt{\mbox{tr}\left(\hat{Z}_{t}^{[2]}\right)}$ $=\left\|\left(\Gamma-\Gamma L_{a_{t-1}}c_{a_{t-1}}^{\top}\right)\cdots\left(\Gamma-\Gamma L_{a_{t-s}}c_{a_{t-s}}^{\top}\right)\hat{z}_{t-s}\right\|_{2}.$ (27)
We know that $\Gamma-\Gamma L_{a_{t}}c_{a_{t}}^{\top}$ is stable for any $c_{A_{t}}\in\mathcal{A}$ based on Theorem 1. Therefore, based on (26) and (27), $\sqrt{\mbox{tr}\left(\hat{Z}_{t}^{[2]}\right)}\succeq\sqrt{\mbox{tr}\left(\hat{Z}_{t}^{[2]}\right)}$ implying that
$\left\|\left(\Gamma-\Gamma L_{a_{t-1}}c_{a_{t-1}}^{\top}\right)\cdots\left(\Gamma-\Gamma L_{a_{t-s}}c_{a_{t-s}}^{\top}\right)\hat{z}_{t-s}\right\|_{2}\leq\sqrt{\mbox{tr}\left(\hat{Z}_{t}\right)}=\left\|\hat{z}_{t}\right\|_{2},$
To bound $\left\|\hat{z}_{t}\right\|_{2}$, assumption 6 states that $\sqrt{\mbox{tr}\left(\mathbb{E}\left[z_{t}z_{t}^{\top}\right]\right)}\leq B_{R}$. Since $z_{t}=\hat{z}_{t}+e_{t}$ and by the orthogonality principle
$\mbox{tr}\left(\mathbb{E}\left[\hat{z}_{t}e_{t}^{\top}+e_{t}\hat{z}_{t}^{\top}\right]\right)=0,$
then using $Z_{t}$ in (22),
$Z_{t}=\mathbb{E}\left[z_{t}z_{t}^{\top}\right]$ $=\mathbb{E}\left[\left(\hat{z}_{t}+e_{t}\right)\left(\hat{z}_{t}+e_{t}\right)^{\top}\mid\mathcal{F}_{t-1}\right]$
$=\mathbb{E}\left[\hat{z}_{t}\hat{z}_{t}^{\top}+\hat{z}_{t}e_{t}^{\top}+e_{t}\hat{z}_{t}^{\top}+e_{t}e_{t}^{\top}\mid\mathcal{F}_{t-1}\right]$
$=\mathbb{E}\left[\hat{z}_{t}\hat{z}_{t}^{\top}\mid\mathcal{F}_{t-1}\right]+\mathbb{E}\left[\hat{z}_{t}e_{t}^{\top}+e_{t}\hat{z}_{t}^{\top}\mid\mathcal{F}_{t-1}\right]+\mathbb{E}\left[e_{t}e_{t}^{\top}\mid\mathcal{F}_{t-1}\right]$
$=\mathbb{E}\left[\hat{z}_{t}\hat{z}_{t}^{\top}\mid\mathcal{F}_{t-1}\right]+\mathbb{E}\left[e_{t}e_{t}^{\top}\mid\mathcal{F}_{t-1}\right]$
$=\hat{Z}_{t}+P_{t},$
where $\hat{Z}_{t}\triangleq\mathbb{E}\left[z_{t}z_{t}^{\top}\mid\mathcal{F}_{t-1}\right]$ and $\hat{P}_{t}\triangleq\mathbb{E}\left[e_{t}e_{t}^{\top}\mid\mathcal{F}_{t-1}\right]$. Since $\hat{Z}_{t}\succeq 0$ and $P_{t}\succeq 0$, then
$\sqrt{\mbox{tr}\left(Z_{t}\right)}=\sqrt{\mbox{tr}\left(\hat{Z}_{t}+P_{t}\right)}\geq\sqrt{\mbox{tr}\left(\hat{Z}_{t}\right)}$
$\Rightarrow\sqrt{\mbox{tr}\left(\hat{Z}_{t}\right)}\leq B_{R}.$
Therefore, using the Markov inequality [25], with a probability of at least $1-\delta$,
$\left|\beta_{A_{t}}\right|\leq\left\|c_{a}\right\|_{2}\frac{\sqrt{\mbox{tr}\left(\mathbb{E}\left[z_{t}z_{t}^{\top}\right]\right)}}{\delta}\leq\left\|c_{a}\right\|_{2}\frac{B_{R}}{\delta}\leq\frac{B_{c}B_{R}}{\delta}.$ (28)
Using inequality (28), inequality (24) can be rewritten as
$\left\|\mathbf{Z}_{\mathcal{T}_{c_{A_{t}}\mid\mathbf{c}}}\mathbf{B}_{\mathcal{T}_{c_{A_{t}}\mid\mathbf{c}}}^{\top}\right\|_{V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)^{-1}}$ $\leq\sum_{i=1}^{N_{a}}\frac{B_{c}B_{R}}{\delta}\left\|\Xi_{t_{i}}\left(\mathbf{c}\right)^{\top}V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)^{-1/2}\right\|_{2}$
$=\sum_{i=1}^{N_{a}}\frac{B_{c}B_{R}}{\delta}\sqrt{\mbox{tr}\left(V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)^{-1}\Xi_{t_{i}}\left(\mathbf{c}\right)\Xi_{t_{i}}\left(\mathbf{c}\right)^{\top}\right)}$
$\leq\sum_{i=1}^{N_{a}}\frac{B_{c}B_{R}}{\delta}\sqrt{\mbox{tr}\left(V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)^{-1}\Xi_{t_{i}}\left(\mathbf{c}\right)\Xi_{t_{i}}\left(\mathbf{c}\right)^{\top}\right)\cdot 1}$
$\overset{(*)}{\leq}\frac{B_{c}B_{R}}{\delta}\sqrt{\sum_{i=1}^{N_{a}}1}\sqrt{\sum_{i=1}^{N_{a}}\mbox{tr}\left(V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)^{-1}\Xi_{t_{i}}\left(\mathbf{c}\right)\Xi_{t_{i}}\left(\mathbf{c}\right)^{\top}\right)}$
$=\sqrt{N_{a}}\frac{B_{c}B_{R}}{\delta}\sqrt{\mbox{tr}\left(V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)^{-1}\sum_{i=1}^{N_{a}}\Xi_{t_{i}}\left(\mathbf{c}\right)\Xi_{t_{i}}\left(\mathbf{c}\right)^{\top}\right)}$
$=\sqrt{N_{a}}\frac{B_{c}B_{R}}{\delta}\sqrt{\mbox{tr}\left(V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)^{-1}\left(V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)-\lambda I\right)\right)},$
$\Rightarrow\left\|\mathbf{Z}_{\mathcal{T}_{c_{A_{t}}\mid\mathbf{c}}}\mathbf{B}_{\mathcal{T}_{c_{A_{t}}\mid\mathbf{c}}}^{\top}\right\|_{V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)^{-1}}\leq\sqrt{N_{a}}\frac{B_{c}B_{R}}{\delta}\sqrt{\mbox{tr}\left(I-\lambda V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)^{-1}\right)},$
where ()(*)( ∗ ) uses the Cauchy-Schwarz inequality. Since $V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)\succeq\lambda I$ based on (11), then the difference $I-\lambda V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)\succeq 0$. Therefore, no further effort is needed to show that inequality (22) is satisfied with a probability of at least $1-\delta$.
Using Lemma 2, Theorem 3 can be proven.
Theorem 3.
Let $\hat{G}_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)$ be identified based on (10) where the reward $\mathbf{X}_{\mathcal{T}_{c_{a}\mid\mathbf{c}}}$ has expression (9). Then, the following inequality is satisfied with a probability of at least $\left(1-\delta\right)^{2}$, where $\delta\in(0,1)$ are designed parameters for tolerated failure probabilities:
$\left\|\hat{G}_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)-G_{c_{a}\mid\mathbf{c}}\right\|_{V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)}\leq e_{c_{a}\mid\mathbf{c}}\left(\delta\right)+b_{c_{a}\mid\mathbf{c}}\left(\delta\right),$ (29)
where $e_{c_{a}\mid\mathbf{c}}\left(\delta\right)$ and $b_{c_{a}\mid\mathbf{c}}\left(\delta\right)$ are defined in (14) and (15), respectively.
Proof.
Let $\mathbf{X}_{\mathcal{T}_{c_{a}\mid\mathbf{c}}}$ be expressed using (9). The norm $\left\|\hat{G}_{c_{a}\mid\mathbf{c}}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)-G_{c_{a}\mid\mathbf{c}}\right\|_{2}$ can be expressed the following way
$\hat{G}_{c_{a}\mid\mathbf{c}}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)-G_{c_{a}\mid\mathbf{c}}=-\lambda G_{c_{a}\mid\mathbf{c}}V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)^{-1}+\left(\mathbf{B}_{\mathcal{T}_{c_{A_{t}}\mid\mathbf{c}}}+\mathbf{E}_{\mathcal{T}_{c_{A_{t}}\mid\mathbf{c}}}\right)\mathbf{Z}_{\mathcal{T}_{c_{A_{t}}\mid\mathbf{c}}}^{\top}V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)^{-1},$
which implies that:
$\left(\hat{G}_{c_{a}\mid\mathbf{c}}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)^{\top}-G_{c_{a}\mid\mathbf{c}}^{\top}\right)V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)$ $=-\lambda G_{c_{a}\mid\mathbf{c}}^{\top}+\left(\mathbf{B}_{\mathcal{T}_{c_{A_{t}}\mid\mathbf{c}}}+\mathbf{E}_{\mathcal{T}_{c_{A_{t}}\mid\mathbf{c}}}\right)\mathbf{Z}_{\mathcal{T}_{c_{A_{t}}\mid\mathbf{c}}}^{\top},$
$\left(\hat{G}_{c_{a}\mid\mathbf{c}}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)-G_{c_{a}\mid\mathbf{c}}\right)^{\top}V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)$ $=\left(-\lambda G_{c_{a}\mid\mathbf{c}}+\mathbf{Z}_{\mathcal{T}_{c_{A_{t}}\mid\mathbf{c}}}\left(\mathbf{B}_{\mathcal{T}_{c_{A_{t}}\mid\mathbf{c}}}+\mathbf{E}_{\mathcal{T}_{c_{A_{t}}\mid\mathbf{c}}}\right)^{\top}\right)^{\top},$
and consequently:
$\left\|\hat{G}_{c_{a}\mid\mathbf{c}}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)-G_{c_{a}\mid\mathbf{c}}\right\|_{V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)}=\left\|-\lambda G_{c_{a}\mid\mathbf{c}}+\mathbf{Z}_{\mathcal{T}_{c_{A_{t}}\mid\mathbf{c}}}\left(\mathbf{B}_{\mathcal{T}_{c_{A_{t}}\mid\mathbf{c}}}+\mathbf{E}_{\mathcal{T}_{c_{A_{t}}\mid\mathbf{c}}}\right)^{\top}\right\|_{V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)^{-1}}.$
Using the triangle inequality, it follows from (C.1) that:
$\left\|\hat{G}_{c_{a}\mid\mathbf{c}}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)-G_{c_{a}\mid\mathbf{c}}\right\|_{V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)}\leq\left\|\lambda G_{c_{a}\mid\mathbf{c}}\right\|_{V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)^{-1}}+\left\|\mathbf{Z}_{\mathcal{T}_{c_{A_{t}}\mid\mathbf{c}}}\mathbf{B}_{\mathcal{T}_{c_{A_{t}}\mid\mathbf{c}}}^{\top}\right\|_{V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)^{-1}}\\ +\left\|\mathbf{Z}_{\mathcal{T}_{c_{A_{t}}\mid\mathbf{c}}}\mathbf{E}_{\mathcal{T}_{c_{A_{t}}\mid\mathbf{c}}}^{\top}\right\|_{V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)^{-1}}.$ (30)
As it is shown in Lemma 2, inequality (22) is satisfied with a probability of at least $1-\delta$. For the term $\varepsilon_{a;\tau}$, since $\varepsilon_{a;\tau}$ is conditionally $B_{R}$-sub-Gaussian on $\mathcal{F}_{t-1}$ and $\Xi_{t}$ is $\mathcal{F}_{t-1}$ measurable, then the conditions described in [17], Theorem 1, are satisfied. Therefore, the following inequality is satisfied with a probability of at least $1-\delta$:
$\left\|\mathbf{Z}_{\mathcal{T}_{c_{A_{t}}\mid\mathbf{c}}}\mathbf{E}_{\mathcal{T}_{c_{A_{t}}\mid\mathbf{c}}}^{\top}\right\|_{V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)^{-1}}\leq\sqrt{2B_{R}^{2}\log\left(\frac{1}{\delta}\frac{\det(V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right))^{1/2}}{\det(\lambda I)^{1/2}}\right)}.$
Finally, using Assumption 5, the term $\lambda G_{c_{a}\mid\mathbf{c}}$ can be bounded using the Cauchy-Schwarz inequality
$\left\|\lambda G_{c_{a}\mid\mathbf{c}}\right\|_{V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)^{-1}}$ $\leq\left|\lambda\right|\sqrt{\mbox{tr}\left(V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)^{-1}\right)}\left\|G_{c_{a}\mid\mathbf{c}}\right\|_{2}$
$\leq\left|\lambda\right|\sqrt{\mbox{tr}\left(V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)^{-1}\right)}B_{G}$
Therefore, the bound (29) is obtained. ∎
Theorem 3 states that with a probability of at least $\left(1-\delta\right)^{2}$, the weighted-norm in (29) is upper-bounded by two terms, (14) and (15). Note that the terms $\delta$ are within the interval $(0,1)$, where the closer $\delta$ is to 00, the larger the terms (14) and (15) are. The opposite occurs if $\delta$ are closer to $1$.
C.2 Prediction Error
Using Theorem 3, we can prove a bound on the prediction error (21).
Theorem 4.
With a probability of at least $\left(1-\delta\right)^{2}$, the following inequality is satisfied
$\hat{G}_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)^{\top}\Xi\left(\mathbf{c}\right)-G_{c_{a}\mid\mathbf{c}}^{\top}\Xi\left(\mathbf{c}\right)\leq\sigma\left(\delta,c_{a}\mid\mathbf{c}\right),$ (31)
where $\sigma\left(\delta,c_{a}\mid\mathbf{c}\right)$ is defined to be
$\sigma\left(\delta,c_{a}\mid\mathbf{c}\right)\triangleq\left(e_{c_{a}\mid\mathbf{c}}\left(\delta\right)+b_{c_{a}\mid\mathbf{c}}\left(\delta\right)\right)\sqrt{\Xi_{t}\left(\mathbf{c}\right)^{\top}V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)^{-1}\Xi_{t}\left(\mathbf{c}\right)}.$ (32)
Proof.
Let there be the following optimization problem
$\begin{array}[]{cc}\underset{a\in[k]}{\arg\max}&\tilde{G}_{c_{a}\mid\mathbf{c}}^{\top}\Xi\left(\mathbf{c}\right)\\ \mbox{s.t.}&\left\|\hat{G}_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)-\tilde{G}_{c_{a}\mid\mathbf{c}}\right\|_{V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)}^{2}\leq\left(e_{c_{a}\mid\mathbf{c}}\left(\delta\right)+b_{c_{a}\mid\mathbf{c}}\left(\delta\right)\right)^{2}\end{array}.$ (33)
To solve (33), we have the following Lagrangian
$L\left(\tilde{G}_{c_{a}\mid\mathbf{c}},\alpha\right)=\tilde{G}_{c_{a}\mid\mathbf{c}}^{\top}\Xi\left(\mathbf{c}\right)-\frac{\alpha}{2}\left(\left\|\hat{G}_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)-\tilde{G}_{c_{a}\mid\mathbf{c}}\right\|_{V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)}^{2}-\left(e_{c_{a}\mid\mathbf{c}}\left(\delta\right)+b_{c_{a}\mid\mathbf{c}}\left(\delta\right)\right)^{2}\right),$
$L\left(\tilde{G}_{c_{a}\mid\mathbf{c}},\alpha\right)=\tilde{G}_{c_{a}\mid\mathbf{c}}^{\top}\Xi\left(\mathbf{c}\right)\\ -\frac{\alpha}{2}\left(\left(\hat{G}_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)-\tilde{G}_{c_{a}\mid\mathbf{c}}\right)^{\top}V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)\left(\hat{G}_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)-\tilde{G}_{c_{a}\mid\mathbf{c}}\right)-\left(e_{c_{a}\mid\mathbf{c}}\left(\delta\right)+b_{c_{a}\mid\mathbf{c}}\left(\delta\right)\right)^{2}\right),$ (34)
$L\left(\tilde{G}_{c_{a}\mid\mathbf{c}},\alpha\right)=\tilde{G}_{c_{a}\mid\mathbf{c}}^{\top}\Xi\left(\mathbf{c}\right)-\frac{\alpha}{2}\hat{G}_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)^{\top}V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)\hat{G}_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)\\ +\alpha\hat{G}_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)^{\top}V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)\tilde{G}_{c_{a}\mid\mathbf{c}}-\frac{\alpha}{2}\tilde{G}_{c_{a}\mid\mathbf{c}}^{\top}V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)\tilde{G}_{c_{a}\mid\mathbf{c}}+\frac{\alpha}{2}\left(e_{c_{a}\mid\mathbf{c}}\left(\delta\right)+b_{c_{a}\mid\mathbf{c}}\left(\delta\right)\right)^{2}.$ (35)
Therefore, the partial derivatives of the Lagrangian are
$\frac{\partial L\left(\tilde{G}_{c_{a}\mid\mathbf{c}},\alpha\right)}{\partial\tilde{G}_{c_{a}\mid\mathbf{c}}}$ $=\Xi\left(\mathbf{c}\right)^{\top}-\alpha\tilde{G}_{c_{a}\mid\mathbf{c}}^{\top}V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)+\alpha\hat{G}_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)^{\top}V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)$
$\frac{\partial L\left(\tilde{G}_{c_{a}\mid\mathbf{c}},\alpha\right)}{\partial\alpha}$ $=-\frac{1}{2}\left(\left\|\hat{G}_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)-\tilde{G}_{c_{a}\mid\mathbf{c}}\right\|_{V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)}^{2}+\left(e_{c_{a}\mid\mathbf{c}}\left(\delta\right)+b_{c_{a}\mid\mathbf{c}}\left(\delta\right)\right)^{2}\right).$
Setting $\partial L\left(\tilde{G}_{c_{a}\mid\mathbf{c}},\alpha\right)/\partial\tilde{G}_{c_{a}\mid\mathbf{c}}=0$ and solving $\tilde{G}_{c_{a}\mid\mathbf{c}}$ provides
$\Xi\left(\mathbf{c}\right)^{\top}-\alpha\tilde{G}_{c_{a}\mid\mathbf{c}}^{\top}V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)+\alpha\hat{G}_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)^{\top}V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)$ $=0$
$\Xi\left(\mathbf{c}\right)^{\top}+\alpha\hat{G}_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)^{\top}V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)$ $=\alpha\tilde{G}_{c_{a}\mid\mathbf{c}}^{\top}V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right),$
$\Rightarrow\tilde{G}_{c_{a}\mid\mathbf{c}}^{\top}=\Xi\left(\mathbf{c}\right)^{\top}\frac{V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)^{-1}}{\alpha}+\hat{G}_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)^{\top}.$ (36)
Plugging in $\tilde{G}_{c_{a}\mid\mathbf{c}}$ in (35) provides
$\frac{1}{2}\left(\left\|\hat{G}_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)-\left(\hat{G}_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)-\frac{V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)^{-1}}{\alpha}\Xi\left(\mathbf{c}\right)\right)\right\|_{V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)}^{2}-\left(e_{c_{a}\mid\mathbf{c}}\left(\delta\right)+b_{c_{a}\mid\mathbf{c}}\left(\delta\right)\right)^{2}\right)$ $=0$
$\frac{1}{2}\left(\left\|\frac{V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)^{-1}}{\alpha}\Xi\left(\mathbf{c}\right)\right\|_{V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)}^{2}-\left(e_{c_{a}\mid\mathbf{c}}\left(\delta\right)+b_{c_{a}\mid\mathbf{c}}\left(\delta\right)\right)^{2}\right)$ $=0$
$\frac{1}{2}\left(\frac{\Xi\left(\mathbf{c}\right)^{\top}V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)^{-1}V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)^{-1}\Xi\left(\mathbf{c}\right)}{\alpha^{2}}-\left(e_{c_{a}\mid\mathbf{c}}\left(\delta\right)+b_{c_{a}\mid\mathbf{c}}\left(\delta\right)\right)^{2}\right)$ $=0$
$\frac{1}{2}\left(\frac{\Xi\left(\mathbf{c}\right)^{\top}V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)^{-1}\Xi\left(\mathbf{c}\right)}{\alpha^{2}}-\left(e_{c_{a}\mid\mathbf{c}}\left(\delta\right)+b_{c_{a}\mid\mathbf{c}}\left(\delta\right)\right)^{2}\right)$ $=0,$
$\frac{\Xi\left(\mathbf{c}\right)^{\top}V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)^{-1}\Xi\left(\mathbf{c}\right)}{\alpha^{2}}$ $=\left(e_{c_{a}\mid\mathbf{c}}\left(\delta\right)+b_{c_{a}\mid\mathbf{c}}\left(\delta\right)\right)^{2}$
$\frac{\Xi\left(\mathbf{c}\right)^{\top}V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)^{-1}\Xi\left(\mathbf{c}\right)}{\left(e_{c_{a}\mid\mathbf{c}}\left(\delta\right)+b_{c_{a}\mid\mathbf{c}}\left(\delta\right)\right)^{2}}$ $=\alpha^{2},$
$\Rightarrow\alpha=\sqrt{\frac{\Xi\left(\mathbf{c}\right)^{\top}V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)^{-1}\Xi\left(\mathbf{c}\right)}{\left(e_{c_{a}\mid\mathbf{c}}\left(\delta\right)+b_{c_{a}\mid\mathbf{c}}\left(\delta\right)\right)^{2}}}.$ (37)
Finally, plugging in (37) and (36) into $\tilde{G}_{c_{a}\mid\mathbf{c}}^{\top}\Xi\left(\mathbf{c}\right)$ provides
$\tilde{G}_{c_{a}\mid\mathbf{c}}^{\top}\Xi\left(\mathbf{c}\right)=\hat{G}_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)^{\top}\Xi\left(\mathbf{c}\right)+\Xi\left(\mathbf{c}\right)^{\top}V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)^{-1}\Xi\left(\mathbf{c}\right)\sqrt{\frac{\left(e_{c_{a}\mid\mathbf{c}}\left(\delta\right)+b_{c_{a}\mid\mathbf{c}}\left(\delta\right)\right)^{2}}{\Xi\left(\mathbf{c}\right)^{\top}V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)^{-1}\Xi\left(\mathbf{c}\right)}}.$
The solution therefore for (33) is
$\hat{G}_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)^{\top}\Xi\left(\mathbf{c}\right)+\left(e_{c_{a}\mid\mathbf{c}}\left(\delta\right)+b_{c_{a}\mid\mathbf{c}}\left(\delta\right)\right)\sqrt{\Xi\left(\mathbf{c}\right)^{\top}V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)^{-1}\Xi\left(\mathbf{c}\right)}.$ (38)
If we define $\sigma\left(\delta,c_{a}\mid\mathbf{c}\right)$ to be (32), then the inequality (31) is satisfied with a probability of at least $\left(1-\delta\right)^{2}$. ∎
The following lemma is provided for proving the bound of (32).
Lemma 3.
The upper bound of $\sigma\left(\delta,c_{a}\mid\mathbf{c}\right)$ is satisfied with a probability of at least $1-\delta$ where the bound is
$\sigma\left(\delta,c_{a}\mid\mathbf{c}\right)\leq B\left(\delta,c_{a_{t}^{*}}\mid\mathbf{c}\right),$ (39)
where $B\left(\delta,c_{a_{t}^{*}}\mid\mathbf{c}\right)$ is defined to be
$B\left(\delta,c_{a_{t}^{*}}\mid\mathbf{c}\right)\triangleq\sqrt{2B_{R}^{2}\log\left(\frac{1}{\delta}\frac{\left(s\lambda+n\frac{\mathbb{E}\left[\left\|\Xi_{t}\left(\mathbf{c}\right)\right\|_{2}\right]}{\delta}\right)^{s/2}}{\lambda^{s/2}}\right)}\sqrt{\frac{s}{\lambda}}\frac{\mathbb{E}\left[\left\|\Xi_{\tau}\left(\mathbf{c}\right)\right\|_{2}\right]}{\delta}\\ +\sqrt{n}B_{c}B_{R}\sqrt{s}\sqrt{\frac{s}{\lambda}}\frac{\mathbb{E}\left[\left\|\Xi_{\tau}\left(\mathbf{c}\right)\right\|_{2}\right]}{\delta}+\lambda B_{G}\sqrt{\frac{s}{\lambda}}\frac{\mathbb{E}\left[\left\|\Xi_{\tau}\left(\mathbf{c}\right)\right\|_{2}\right]}{\delta}.$ (40)
Proof.
The upper bound of $\sigma\left(\delta,c_{a}\mid\mathbf{c}\right)$ is as follows: given that the following inequality is satisfied with a probability of least $1-\delta$:
$\left\|\Xi_{t}\left(\mathbf{c}\right)\right\|_{2}\leq\frac{\mathbb{E}\left[\left\|\Xi_{t}\left(\mathbf{c}\right)\right\|_{2}\right]}{\delta},$ (41)
then according to Lemma 11 in [17], $e_{c_{a}\mid\mathbf{c}}\left(\delta\right)$ is upper bounded as (with a probability of at least $1-\delta$)
$e_{c_{a}\mid\mathbf{c}}\left(\delta\right)$ $\leq\sqrt{2B_{R}^{2}\log\left(\frac{1}{\delta}\frac{\left(\mbox{tr}\left(\lambda I_{s}\right)+\left|\mathcal{T}_{c_{a}\mid\mathbf{c}}\right|\frac{\mathbb{E}\left[\left\|\Xi_{t}\left(\mathbf{c}\right)\right\|_{2}\right]}{\delta}\right)^{s/2}}{\sqrt{\det\left(\lambda I_{s}\right)}}\right)}$
$\leq\sqrt{2B_{R}^{2}\log\left(\frac{1}{\delta}\frac{\left(s\lambda+n\frac{\mathbb{E}\left[\left\|\Xi_{t}\left(\mathbf{c}\right)\right\|_{2}\right]}{\delta}\right)^{s/2}}{\lambda^{s/2}}\right)}.$
The term $b_{c_{a}\mid\mathbf{c}}\left(\delta\right)$ is based on bound (22). To bound $\sqrt{\mbox{tr}\left(I_{s}-\lambda V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)^{-1}\right)}$ in (22), note that since $V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)\succeq\lambda I$, then
$\sqrt{\mbox{tr}\left(I_{s}-\lambda V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)^{-1}\right)}\leq\sqrt{\mbox{tr}\left(I_{s}\right)}=\sqrt{s}.$
Therefore, we have the upper bound of $b_{c_{a}\mid\mathbf{c}}\left(\delta\right)$:
$b_{c_{a}\mid\mathbf{c}}\left(\delta\right)\leq\sqrt{n}B_{c}B_{R}\sqrt{s}.$
Finally, the term $\sqrt{\Xi_{\tau}\left(\mathbf{c}\right)^{\top}V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)^{-1}\Xi_{\tau}\left(\mathbf{c}\right)}$ has the following bound with a probability of at least $1-\delta$
$\sqrt{\Xi_{\tau}\left(\mathbf{c}\right)^{\top}V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)^{-1}\Xi_{\tau}\left(\mathbf{c}\right)}$ $\leq\sqrt{\mbox{tr}\left(V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)^{-1}\right)}\left\|\Xi_{\tau}\left(\mathbf{c}\right)\right\|_{2}$
$\leq\sqrt{\mbox{tr}\left(\frac{I_{s}}{\lambda}\right)}\left\|\Xi_{\tau}\left(\mathbf{c}\right)\right\|_{2},$
$\Rightarrow\sqrt{\Xi_{\tau}\left(\mathbf{c}\right)^{\top}V_{a}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)^{-1}\Xi_{\tau}\left(\mathbf{c}\right)}\leq\sqrt{\frac{s}{\lambda}}\frac{\mathbb{E}\left[\left\|\Xi_{\tau}\left(\mathbf{c}\right)\right\|_{2}\right]}{\delta}.$ (42)
where the final inequality (42) uses (41). Therefore, the bound for $\sigma\left(\delta,c_{a}\mid\mathbf{c}\right)$ is (39), which satisfied with a probability of at least $1-\delta$. ∎
C.3 Regret Performance
The following theorem provides a bound for regret.
Theorem 5.
Regret satisfies the following inequality with a probability of at least $\left(1-\delta\right)^{4}$
$R_{n}\leq\sum_{a=1}^{k}2nB_{c}^{2}B_{R}^{2}\left(1-\left(1-\delta\right)^{4}\left(1-\exp\left(\frac{-\phi_{a,a_{t}^{*}}^{2}}{2\Delta G_{c_{a}\mid\mathbf{c}}^{\top}\Sigma_{\Xi_{t}\left(\mathbf{c}\right)}\Delta G_{c_{a}\mid\mathbf{c}}}\right)\right)\right),$ (43)
where $a_{t}^{*}\in[k]$ is the index of the optimal action $c_{a_{t}^{*}}\in\mathcal{A}$ at round $t$. Terms $\Delta G_{c_{a}\mid\mathbf{c}}$, $\phi_{a,a_{t}^{*}}^{2}$, and $\Sigma_{\Xi_{t}\left(\mathbf{c}\right)}$ are defined to be
$\Delta G_{c_{a}\mid\mathbf{c}}$ $\triangleq G_{c_{a^{*}}\mid\mathbf{c}}-G_{c_{a}\mid\mathbf{c}}$ (44)
$\phi_{a,a^{*}}^{2}$ $\triangleq u_{a^{*}}-u_{a}+\sigma\left(\delta,c_{a^{*}}\mid\mathbf{c}\right)+\sigma\left(\delta,c_{a}\mid\mathbf{c}\right)$ (45)
$\Sigma_{\Xi_{t}\left(\mathbf{c}\right)}$ $\triangleq\mathbb{E}\left[\Xi_{t}\left(\mathbf{c}\right)\Xi_{t}\left(\mathbf{c}\right)^{\top}\right],$ (46)
and $\sigma\left(\delta,c_{a}\mid\mathbf{c}\right)$ is defined to be (32).
Proof.
Define $c_{a^{*}}\in\mathcal{A}$ as the optimal action, i.e. choosing action $c_{a^{*}}\in\mathcal{A}$ samples a reward $X_{t}^{*}$ from (1) such that for any action $c_{a}\in\mathcal{A}$ with a sampled reward $X_{t}$, the reward $X_{t}^{*}\geq X_{t}$. The probability of choosing the optimal action $c_{a^{*}}\in\mathcal{A}$ is based on the following event
$\hat{G}_{c_{a^{*}}\mid\mathbf{c}}\left(\mathcal{T}_{c_{a^{*}}\mid\mathbf{c}}\right)^{\top}\Xi\left(\mathbf{c}\right)+u_{a^{*}}\geq\hat{G}_{c_{a}\mid\mathbf{c}}\left(\mathcal{T}_{c_{a}\mid\mathbf{c}}\right)^{\top}\Xi\left(\mathbf{c}\right)+u_{a}.$
With a probability of at least $\left(1-\delta\right)^{4}$, the following inequality is satisfied
$G_{c_{a^{*}}\mid\mathbf{c}}^{\top}\Xi\left(\mathbf{c}\right)+u_{a^{*}}+\sigma\left(\delta,c_{a^{*}}\mid\mathbf{c}\right)\geq G_{c_{a}\mid\mathbf{c}}^{\top}\Xi\left(\mathbf{c}\right)+u_{a}-\sigma\left(\delta,c_{a}\mid\mathbf{c}\right).$
Rearranging provides the following
$\left(G_{c_{a^{*}}\mid\mathbf{c}}-G_{c_{a}\mid\mathbf{c}}\right)^{\top}\Xi\left(\mathbf{c}\right)\geq u_{a}-\sigma\left(\delta,c_{a}\mid\mathbf{c}\right)-u_{a^{*}}-\sigma\left(\delta,c_{a^{*}}\mid\mathbf{c}\right).$
Define the left side of the inequality as (44) and the right side of inequality (C.3) as (45). Since $\Delta G_{c_{a}\mid\mathbf{c}}^{\top}\left(\Xi\left(\mathbf{c}\right)-\mathbb{E}\left[\Xi\left(\mathbf{c}\right)\right]\right)$ is sub-Gaussian, then using the Cramer-Chernoff bound (based on a Corollary 5.5 in [21]) provides
$\mathbb{P}\left(\Delta G_{c_{a}\mid\mathbf{c}}^{\top}\Xi\left(\mathbf{c}\right)\geq\Delta G_{c_{a}\mid\mathbf{c}}^{\top}\mathbb{E}\left[\Xi\left(\mathbf{c}\right)\right]-\phi_{a,a^{*}}\right)\geq 1-\exp\left(\frac{-\phi_{a,a^{*}}^{2}}{2\Delta G_{c_{a}\mid\mathbf{c}}^{\top}\Sigma_{\Xi_{t}\left(\mathbf{c}\right)}\Delta G_{c_{a}\mid\mathbf{c}}}\right)$ (47)
the optimal action is chosen. The probability of choosing the sub-optimal action is the probability of not choosing the optimal action. This implies that with a probability of at most
$p_{a}\triangleq 1-\left(1-\delta\right)^{4}\left(1-\exp\left(\frac{-\phi_{a,a^{*}}^{2}}{2\Delta G_{c_{a}\mid\mathbf{c}}^{\top}\Sigma_{\Xi_{t}\left(\mathbf{c}\right)}\Delta G_{c_{a}\mid\mathbf{c}}}\right)\right),$
the sub-optimal action is chosen. Using the law of iterated expectations [26], the upper bound for instantaneous regret $\mathbb{E}\left[X_{t}^{*}-X_{t}\right]$ is
$\mathbb{E}\left[X_{t}^{*}-X_{t}\right]=\sum_{a=1}^{k}\mathbb{E}\left[\left\langle\Delta c_{a},z_{t}\right\rangle\mid a\right]\mathbb{P}\left[a\right]\leq\sum_{a=1}^{k}\mathbb{E}\left[\left\langle\Delta c_{a},z_{t}\right\rangle\mid a\right]p_{a},$
$\Rightarrow\mathbb{E}\left[X_{t}^{*}-X_{t}\right]\leq\sum_{a=1}^{k}2B_{c}^{2}B_{R}^{2}\left(1-\left(1-\delta\right)^{4}\left(1-\exp\left(\frac{-\phi_{a,a^{*}}^{2}}{2\Delta G_{c_{a}\mid\mathbf{c}}^{\top}\Sigma_{\Xi_{t}\left(\mathbf{c}\right)}\Delta G_{c_{a}\mid\mathbf{c}}}\right)\right)\right).$
Therefore, regret is upper-bounded by inequality (43). ∎
The regret bound is a function of prediction errors $\sigma\left(\delta,c_{a}\mid\mathbf{c}\right)$ and $\sigma\left(\delta,c_{a^{*}}\mid\mathbf{c}\right)$. To control the impact of prediction error $\sigma\left(\delta,c_{a}\mid\mathbf{c}\right)$, we set $u_{a}$ and $u_{a^{*}}$ to be $u_{a}=\sigma\left(\delta,c_{a}\mid\mathbf{c}\right)$ and $u_{a^{*}}=\sigma\left(\delta,c_{a^{*}}\mid\mathbf{c}\right)$. Therefore, by setting $u_{a}=\sigma\left(\delta,c_{a}\mid\mathbf{c}\right)$ for all $c_{a}\in\mathcal{A}$ and using Theorem 5, Theorem 2 can be easily proven.